McGilla Golf is evaluating a new line of golf clubs. The clubs will sell for $1,080 per set and have a variable cost of $490 per set. The company has spent $177,500 for a marketing study that determined the company will sell 54,500 sets per year for seven years. The marketing study also determined that the company will lose sales of 10,300 sets of its high-priced clubs. The high-priced clubs sell at $1,580 and have variable costs of $710. The company also will increase sales of its cheap clubs by 12,900 sets. The cheap clubs sell for $490 and have variable costs of $220 per set. The fixed costs each year will be $10,050,000. The company has also spent $1,375,000 on research and development for the new clubs. The plant and equipment required will cost $33,950,000 and will be depreciated on a straight-line basis to a zero salvage value. The new clubs also will require an increase in net working capital of $2,770,000 that will be returned at the end of the project. The tax rate is 25 percent and the cost of capital is 15 percent.

    

Suppose you feel that the values are accurate to within only ±10 percent. What are the best-case and worst-case NPVs? (Hint: The price and variable costs for the two existing sets of clubs are known with certainty; only the sales gained or lost are uncertain.) (A negative amount should be indicated by a minus sign. Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
Answer #1

Please find below table useful to compute desired results: -

C D G H Base Sales =B5 Price =B2 Varibale cost =B3 Fixed costs =B12 Sales Lost(Expensive) =B6 Sales Gained(Cheap) =B9 А B 1 PB C D Year Calculation of NPV PV Factor CFAT Best Case Worst Case =1/(1+B17) E31 =-B14+B15) =-(B14+B15) =1/(1+B17) E32 =H26 =

End results would be as follows: -

A E Base Sales Price Varibale cost Fixed costs Sales Lost(Expensive) Sales Gained(Cheap) 54,500 1,080 490 10.050.000 10,300 1B C D G H Year PV Factor 1.0000 0.8696 0.7561 0.6575 0.5718 0.4972 0.4323 0.3759 NPV Calculation of NPV CFAT PV of CFAT Best